OpenStudy (anonymous):

(tan(theta)-1)(cos(theta)+1)=0 on interval 0<=theta<=2pi

OpenStudy (anonymous):

\[(\tan(\theta)-1)(\cos(\theta)+1)=0\]

OpenStudy (aum):

\[ (\tan(\theta)-1) = 0 ~~\text{ or } ~~ (\cos(\theta)+1)=0 \]

OpenStudy (midhun.madhu1987):

a * b = 0 means either a = 0 or b = 0 similarly